OpenStudy (anonymous):

If f(x) = x − 4 and g(x)1/2 = x − 2, find f[g(−6)]. My answer is 9 Am I right?

OpenStudy (mathstudent55):

What does g(x)1/2 = x - 2 mean? What is that "1/2" doing there?

OpenStudy (anonymous):

Oops sorry I wrote it wrong! Its supposed to be If f(x) = x − 4 and g(x)= 1/2x − 2, find f[g(−6)].

OpenStudy (mathstudent55):

Then g(-6) = (1/2)(-6) - 2 = -5 f(-5) = -5 - 4 = -9

OpenStudy (anonymous):

Thanks! And sorry about that! :)

OpenStudy (mathstudent55):

You're welcome and no need to apologize. The important thing is did you understand it?